OpenStudy (lilday12):

Multiply the complex numbers. (6 + 4i)(5 - 9i)

OpenStudy (mathmale):

Be sure you're familiar with the following: real times real number produces a real product. Thus, 6*5 = 30. Real times imaginary => imaginary product. Imaginary quantity times imaginary quantity: The product will be -1. Example: 3i * 2i=-6

OpenStudy (mathmale):

Following these rules, find the product (6 + 4i)(5 - 9i). Please show your work.

OpenStudy (anonymous):

6(5) + 6(-9i) +4i(5) + 4i(-9i) using the foil method which gives you = 30 + -54i + 20i + -36i^2 now remove the i^2: -36i^2 = 36 since i^2 = -1 you just multiply -36(-1) giving you 36 leaving you with: 30 + 36 -54i + 20i So you end up with -34i + 30 + 36 = (66 - 34i) <--- final answer

OpenStudy (anonymous):

best response please.

OpenStudy (anonymous):

He confirmation my work is correct http://www.wolframalpha.com/input/?i=%286+%2B+4i%29%285+-+9i%29

OpenStudy (mathmale):

@lugerm: your obvious know-how is important and is appreciated. However, OpenStudy is a teaching site whereby our focus is to help others in the process of solving problems themselves. Unfortunately, you and @lilday12 were not involved in any two-way dialogue at all; rather, you did the work for him / her and then requested a medal for having done so. This is not teaching worthy of a medal.

OpenStudy (anonymous):

i just started today so can i do the work then ask them to work out the answer?

OpenStudy (anonymous):

@mathmale